OpenStudy (anonymous):

y-y1=m(x-x1) what is y1 and what x1?

pooja195 (pooja195):

The coordinate order (x1,y1) (x2,y2)

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

but like, is xi the y-intercept or the x-intercept?

pooja195 (pooja195):

Its the coordinates given

OpenStudy (anonymous):

ohhhh right okay, just checking. i havent gotten this yet. lol thanks :)

pooja195 (pooja195):

Hmm do you want to do a sample problem?

OpenStudy (anonymous):

i think im okay, i just need to answer what they were.

OpenStudy (smilie2015):

for example u r given (2,1) and m is 5 y-1=5(x-2)

pooja195 (pooja195):

ok

OpenStudy (anonymous):

yes yes thanks :)